# A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare

### v1

@@ Appendix A Comparison of Datasets
MedVQA Curated domain-specific dataset A collection of medical visual question answering pairs, designed to train and evaluate models that interpret medical images and answer related questions. Visual question answering, medical image understanding. Explainability
@@
MedExQA Curated domain-specific dataset A dataset focused on medical examination questions and answers, intended to aid in the development of AI models for medical exam preparation and assessment. Question answering, educational assessment. Explainability
@@
MedMCQA Curated domain-specific dataset A multiple-choice question-answering dataset in the medical domain, aimed at training models to handle medical examinations and practice questions. Multiple-choice question answering, medical education. Explainability
@@
TCM Medical Licensing Examination(MLE) Curated domain-specific dataset A dataset comprising questions and answers from Traditional Chinese Medicine licensing examinations. Educational assessment, question answering. Explainability
@@
Pneumonia Dataset Curated domain-specific dataset Medical images (such as chest X-rays) labeled for the presence or absence of pneumonia, used for training diagnostic models. Image classification, disease detection. Explainability
@@
Montgomery Dataset Curated domain-specific dataset Chest X-ray images with manual segmentations of the lung fields, useful for pulmonary research. Image segmentation, tuberculosis detection. Explainability
@@
Shenzhen Dataset Curated domain-specific dataset Chest X-ray images collected in Shenzhen, China, with annotations for tuberculosis manifestations. Disease classification, image analysis. Explainability
@@
IDRID Dataset Curated domain-specific dataset Retinal images with annotations for diabetic retinopathy lesions, intended for retinal image analysis. Image segmentation, disease grading. Explainability
BrainTumor Dataset Curated domain-specific dataset MRI images of brain tumors with corresponding labels, used for developing diagnostic and segmentation models. Tumor detection, image segmentation. Explainability
@@
Appendix B Comparison of Models
@@
We systematically gathered and analyzed 78 models relevant to studies on trust in LLMs for healthcare. Table 2 provides a comprehensive summary of the LLMs evaluated in these studies, detailing key aspects such as model name, release year, and the institution responsible for its development. Additionally, it specifies the primary task each model is designed for, including medical question-answering, clinical decision support, and biomedical text summarization. To further assess their reliability, we categorize the models based on the dimensions of trustworthiness they address, such as truthfulness, privacy, safety, robustness, fairness and bias, and explainability. This structured overview offers valuable insights into how different LLMs are designed and evaluated to enhance trust in healthcare AI applications.
@@
Models Release Year Institution Primary Task Dimensions
@@
SciBERT 2019 Allen Institute for AI Pre-trained language model specialized for scientific text, particularly biomedical and computer science literature. Fairness and Bias
@@
PaLM-2 2023 Google Multilingual language understanding and generation, with a focus on reasoning and coding tasks. Fairness and Bias
@@
Mixtral-8x70B 2023 Mistral AI Ensemble of language models aimed at improving performance across diverse language tasks. Fairness and Bias, Safety
@@
Med-PaLM 2023 Google Health Specializing in healthcare-related question answering, clinical diagnosis support, and medical literature interpretation. Fairness and Bias
@@
Med-PaLM 2 2024 Google Health Updated version of Med-PaLM, further improving healthcare-related tasks with enhanced accuracy and reliability in medical information retrieval, clinical reasoning, and decision support. Fairness and Bias
@@
Llama-13B 2023 Meta Designed for natural language understanding and generation tasks, such as text summarization, machine translation, and conversational AI. Fairness and Bias
@@
XLNet 2019 Google Research It is used for text classification, question answering, and language modeling tasks. Fairness and Bias
@@
DeBERTa 2020 Microsoft Research Improves BERT and RoBERTa by enhancing the attention mechanism. It performs well in a variety of NLP tasks, such as sentence classification, question answering, and named entity recognition. Fairness and Bias
@@
Llama-7B 2023 Meta Focused on general-purpose natural language understanding and generation, with potential fine-tuning for specific domains like medicine, law, and technology. Fairness and Bias, Truthfulness
@@
Llama 2 70Bchat 2023 Meta Platforms Open-source conversational AI model designed for dialogue and instruction-following tasks. Fairness and Bias, Truthfulness, Safety, Robustness,
@@
GPT-3.5 2022 OpenAI Enhanced language processing capabilities, building upon GPT-3. Fairness and Bias, Truthfulness, Safety, Robustness, Privacy
@@
GPT2 2019 OpenAI Text generation Fairness and Bias, Robustness
PMC Llama 13B 2023 Allen Institute for AI Specialized in medical literature understanding and generation. Fairness and Bias, Robustness
@@
Models Release Year Institution Primary Task Dimensions
@@
GPT-4 2023 OpenAI Advanced language generation and understanding across various domains. Fairness and Bias, Safety, Robustness, Explainability, Privacy
@@
BERT 2018 Google AI Language Pre-trained Transformer model for a wide range of NLP tasks, such as text classification, NER, QA, etc. Fairness and Bias, Safety, Robustness, Truthfulness
@@
LLAMA 2 CHAT 2023 Meta AI Language modeling Robustness, Explainability
@@
MEDALPACA (7B) 2023 medalpaca Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Robustness, Privacy
@@
CLINICAL CAMEL (13B) 2023 the AI and healthcare community Fine-tuned for clinical applications. It is designed to assist with tasks like medical text classification, clinical decision support, information extraction from medical records, and answering clinical questions. Robustness
@@
GPT-2 XL 2019 OpenAI Large-scale language model for text generation and understanding. Robustness
@@
T5-Large 2020 Google Research It treats all NLP tasks as text-to-text tasks, meaning both the input and output are in the form of text, and it’s used for tasks like translation, summarization, and question answering. Robustness
@@
claude-3.5-sonnet 2024 Anthropic It is a variant of Claude, specialized in tasks such as conversational AI, creative writing, poetry generation, and other text-based applications. Robustness
@@
OpenBioLLM-70B 2024 OpenBioAI It is designed to handle tasks such as biological information extraction, gene sequence analysis, protein folding predictions, and other bioinformatics applications. Robustness
@@
BioMistral-7B 2023 Mistral AI Focused on biomedical and healthcare-related text. Its tasks include medical question answering, clinical document analysis, and medical text summarization. Robustness
@@
Medllama3-v20 2024 MedAI Labs Designed to assist in healthcare tasks like clinical reasoning, medical question answering, and patient record analysis. Robustness
ASCLEPIUS (7B) 2023 Asclepius AI Developed for clinical and medical applications, specializing in tasks like diagnosing medical conditions from symptoms, medical text summarization, and extracting structured information from clinical documents. Robustness, Explainability
ALPACA (7B) 2023 Stanford University Fine-tuned version of the LLaMA model aimed at providing high-quality responses to questions, with an emphasis on maintaining ethical and accurate conversational capabilities in diverse domains. Robustness
@@
Models Release Year Institution Primary Task Dimensions
@@
Google’s Bard 2023 Google Conversational AI tool, focused on providing detailed, accurate, and creative responses to user queries. It can handle a variety of tasks, including web search, content generation, and complex QA. Robustness
@@
Text- Davinci-003 2022 OpenAI It is an advanced variant of GPT-3. It is designed for a wide range of natural language understanding and generation tasks, such as answering questions, summarizing text, creative writing, translation, and code generation. Robustness, Truthfulness
@@
LLaMa 2-7B 2023 Meta (formerly Facebook AI Research) Designed to be a general-purpose AI for a wide range of tasks such as text generation, question answering, and summarization, with specific fine-tuning for medical and technical domains. Robustness, Truthfulness, Privacy
@@
ChatGPT 2022 OpenAI Conversational AI Robustness, Truthfulness, Explainability, Privacy
@@
Llama-3.1 2024 Meta AI Multilingual large language model designed for a variety of natural language processing tasks. Safety
@@
ClinicalCamel-70b 2023 the AI and healthcare community Medical language model designed for clinical research applications. Safety, Explainability
@@
Med42-70b 2023 M42 Health Clinical large language model providing high-quality answers to medical questions. Safety, Explainability
@@
GPT-4o 2024 OpenAI Multimodal large language model capable of processing and generating text, audio, and images in real time. Safety, Privacy, Explainability
@@
Mistral 2023 Mistral AI Language model optimized for code generation and reasoning tasks. Safety, Robustness, Explainability
@@
Meditron (7) (70b) 2023 École Polytechnique Fédérale de Lausanne (EPFL) Medical language model fine-tuned for clinical decision support and medical reasoning. Safety, Robustness, Explainability
Claude-2.1 2023 Anthropic General-purpose language model for a wide range of natural language understanding and generation tasks. Safety, Robustness
GPT-J 2021 EleutherAI Open-source language model for text generation and understanding. Safety, Robustness
Vicuna 2023 UC Berkeley and Microsoft Research Conversational AI Safety, Robustness, Truthfulness
@@
Models Release Year Institution Primary Task Dimensions
@@
Medalpaca-13b 2023 medalpaca Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Safety, Truthfulness, Privacy
@@
GPT-3 2020 OpenAI Natural language understanding and generation Truthfulness, Explainability
@@
ALBERT 2019 Google Research Lighter version of BERT that reduces parameters for efficiency while maintaining performance. It excels in tasks such as text classification, named entity recognition, and question answering. Truthfulness
@@
RoBERTa 2019 Facebook AI Research Optimized variant of BERT that removes the Next Sentence Prediction task and trains with more data and for longer periods. It is used for tasks like question answering, sentiment analysis, and text classification. Truthfulness
@@
BlueBERT 2019 NIH and Stanford University BERT-based model pre-trained on clinical and biomedical text. It is designed for healthcare-related tasks, including clinical text classification, named entity recognition, and medical question answering. Truthfulness
@@
ClinicalBERT 2019 University of Pennsylvania Variant of BERT fine-tuned on clinical texts, tailored for clinical NLP tasks like named entity recognition, clinical event extraction, and question answering in the medical domain. Truthfulness
@@
TAPAS 2020 Google Research Designed for answering questions based on tabular data. It is used for tasks like extracting structured information from tables and processing queries in tabular datasets. Truthfulness
@@
LLaMA-2 13B 2023 Meta Advanced variant of Meta’s LLaMA series, designed for text generation, question answering, summarization, and other NLP tasks. Truthfulness, Explainability, Privacy
@@
MPT 2023 MosaicML General-purpose LLM for text generation, summarization, language understanding, and reasoning tasks. Fine-tuned for downstream applications such as chatbot development, code generation, and other NLP tasks. Truthfulness
BLIP2 2023 Salesforce Bootstrapping language-image pre-training, designed to bridge vision-language models with large language models for improved visual understanding and generation. Truthfulness
InstructBLIP-7b/13b 2023 Salesforce Visual instruction-tuned versions of BLIP-2, utilizing Vicuna-7B and Vicuna-13B language models, respectively, to enhance vision-language understanding through instruction tuning. Truthfulness
LLaVA1.5-7b/13b 2023 Microsoft Large language and vision assistant models with 7B and 13B parameters, respectively, designed for multimodal tasks by integrating visual information into language models. Truthfulness
mPLUGOwl2 2023 Zhejiang University Multimodal pre-trained language model designed to handle various vision-language tasks, including image captioning and visual question answering. Truthfulness
@@
Models Release Year Institution Primary Task Dimensions
@@
XrayGPT 2023 University of Toronto Specialized model for generating radiology reports from chest X-ray images, aiming to assist in medical image interpretation. Truthfulness
@@
MiniGPT4 2023 King Abdullah University of Science and Technology A lightweight multimodal model designed to align vision and language models efficiently, facilitating tasks like image captioning and visual question answering. Truthfulness
@@
RadFM 2023 Stanford University Foundation model tailored for radiology, focusing on interpreting medical images and integrating findings with clinical language models. Truthfulness
@@
Alpaca-LoRA 2023 Stanford University It focuses on achieving good performance in tasks such as question answering and personalized dialogue. Truthfulness
@@
Robin- medical 2023 Robin Health Fine-tuned for medical applications, including clinical decision support, medical question answering, and health record analysis. Truthfulness
@@
Flan-T5 2021 Google Research Optimized for tasks like question answering, text summarization, and sentence classification, across a variety of domains. Truthfulness
@@
BioBERT 2019 Korea University Biomedical language representation learning, enhancing performance on tasks like named entity recognition, relation extraction, and question answering within the biomedical domain. Truthfulness
@@
Falcon Instruct (7B and 40B) 2023 Technology Innovation Institute (TII), UAE. Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
Mistral Instruct (7B) 2023 Mistral AI Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
Falcon 2023 Technology Innovation Institute (TII), UAE. General-purpose language model optimized for text understanding, generation, question answering, and reasoning tasks. Focused on efficient deployment for industry-scale applications. Truthfulness, Robustness
LLaVA-Med 2024 Microsoft Large language and vision assistant for biomedicine, trained to handle visual instruction tasks in the biomedical field, aiming for capabilities similar to GPT-4. Truthfulness, Explainability
GPT-4o-mini 2024 OpenAI Natural language processing (NLP), text generation, and understanding. Explainability
ASCLEPIUS (13B) 2023 Asclepius AI Medical NLP, clinical text analysis, and healthcare-related tasks. Explainability
@@
Models Release Year Institution Primary Task Dimensions
@@
MedViLaM 2023 Xu et al. (2024) Medical vision-language tasks, combining image and text analysis for healthcare. Explainability
@@
Med-MoE 2023 Jiang et al. (2024) Medical NLP, leveraging Mixture of Experts (MoE) for specialized healthcare tasks. Explainability
@@
Gemini Pro 2023 Google DeepMind Multi-modal NLP, combining text, image, and other data types for advanced AI tasks Explainability
@@
AlpaCare (7B) (13B) 2023 Zhang et al. (2023) Healthcare-focused NLP, clinical text analysis, and medical decision support Explainability
@@
Yi (6B) 2023 01.AI (China) General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
@@
Phi-2 (2.7B) 2023 Microsoft Lightweight NLP, text generation, and fine-tuning for specific tasks. Explainability

### v2

@@ Appendix B Comparison of Datasets
MedVQA Curated domain-specific dataset 794 image-question-answer triplets. A collection of medical visual question answering pairs, designed to train and evaluate models that interpret medical images and answer related questions. (Med-QA) Visual question answering, medical image understanding. Explainability
@@
MedExQA Curated domain-specific dataset 965 multiple-choice medical questions. A dataset focused on medical examination questions and answers, intended to aid in the development of AI models for medical exam preparation and assessment. (Med-QA) Question answering, educational assessment. Explainability
@@
MedMCQA Curated domain-specific dataset 194,000 multiple-choice questions from AIIMS and NEET PG entrance exams, covering 2,400 healthcare topics across 21 medical subjects. A multiple-choice question-answering dataset in the medical domain, aimed at training models to handle medical examinations and practice questions. (Med-QA) Multiple-choice question answering, medical education. Explainability
@@
TCM Medical Licensing Examination(MLE) Curated domain-specific dataset 600 multiple-choice questions. A dataset comprising questions and answers from Traditional Chinese Medicine licensing examinations. (Med-QA) Educational assessment, question answering. Explainability
Datasets Data Type Content Task Dimensions
@@
Pneumonia Dataset Curated domain-specific dataset 5,863 images. Medical images (such as chest X-rays) labeled for the presence or absence of pneumonia, used for training diagnostic models. (Med-IE) Image classification, disease detection. Explainability
@@
Montgomery Dataset Curated domain-specific dataset X-ray Set comprises 138 posterior-anterior chest X-ray images, with 80 normal and 58 abnormal cases indicative of tuberculosis. Chest X-ray images with manual segmentations of the lung fields, useful for pulmonary research. (Med-IE) Image segmentation, tuberculosis detection. Explainability
@@
Shenzhen Dataset Curated domain-specific dataset Chest X-ray dataset comprises 662 frontal chest X-rays, including 326 normal cases and 336 cases with manifestations of tuberculosis. Chest X-ray images collected in Shenzhen, China, with annotations for tuberculosis manifestations. (Med-IE) Disease classification, image analysis. Explainability
@@
IDRID Dataset Curated domain-specific dataset 1,113 images. Retinal images with annotations for diabetic retinopathy lesions, intended for retinal image analysis. (Med-IE) Image segmentation, disease grading. Explainability
MIMIC IV Curated Real-World Clinical Dataset Over 300,000 hospital admissions from Beth Israel Deaconess Medical Center covering de-identified EHR data including demographics, vital signs, medications, diagnoses, and clinical notes (Med-IE / Med-QA / Med-Gen) Used for tasks such as medical code prediction, patient outcome forecasting, clinical summarization, and question answering Explainability
@@
Appendix C Comparison of Models
@@
We systematically gathered and analyzed 81 models relevant to studies on trust in LLMs for healthcare. Table 2 provides a comprehensive summary of the LLMs evaluated in these studies, detailing key aspects such as model name, release year, openness, architecture, and the institution responsible for its development. Additionally, it specifies the primary task each model is designed for, including medical question-answering, clinical decision support, and biomedical text summarization. To further assess their reliability, we categorize the models based on the dimensions of trustworthiness they address, such as truthfulness, privacy, safety, robustness, fairness and bias, and explainability. This structured overview offers valuable insights into how different LLMs are designed and evaluated to enhance trust in healthcare AI applications.
@@
Models Release Year Institution Openness Architecture Primary Task Dimensions
@@
SciBERT 2019 Allen Institute for AI Open-source Encoder-only Pre-trained language model specialized for scientific text, particularly biomedical and computer science literature. Fairness and Bias
@@
PaLM-2 2023 Google Closed-source Decoder-only Multilingual language understanding and generation, with a focus on reasoning and coding tasks. Fairness and Bias
@@
Mixtral-8x70B 2023 Mistral AI Open-source Decoder-only Ensemble of language models aimed at improving performance across diverse language tasks. Fairness and Bias, Safety
@@
Med-PaLM 2023 Google Health Closed-source Decoder-only Specializing in healthcare-related question answering, clinical diagnosis support, and medical literature interpretation. Fairness and Bias
@@
Med-PaLM 2 2024 Google Health Closed-source Encoder-decoder Updated version of Med-PaLM, further improving healthcare-related tasks with enhanced accuracy and reliability in medical information retrieval, clinical reasoning, and decision support. Fairness and Bias
@@
Llama-13B 2023 Meta Open-source Decoder-only Designed for natural language understanding and generation tasks, such as text summarization, machine translation, and conversational AI. Fairness and Bias
@@
XLNet 2019 Google Research Open-source Encoder-only It is used for text classification, question answering, and language modeling tasks. Fairness and Bias
@@
DeBERTa 2020 Microsoft Research Open-source Encoder-only Improves BERT and RoBERTa by enhancing the attention mechanism. It performs well in a variety of NLP tasks, such as sentence classification, question answering, and named entity recognition. Fairness and Bias
@@
Llama-7B 2023 Meta Open-source Decoder-only Focused on general-purpose natural language understanding and generation, with potential fine-tuning for specific domains like medicine, law, and technology. Fairness and Bias, Truthfulness
@@
Llama 2 70Bchat 2023 Meta Platforms Open-source Decoder-only Open-source conversational AI model designed for dialogue and instruction-following tasks. Fairness and Bias, Truthfulness, Safety, Robustness,
@@
GPT-3.5 2022 OpenAI Closed-source Decoder-only Enhanced language processing capabilities, building upon GPT-3. Fairness and Bias, Truthfulness, Safety, Robustness, Privacy
@@
GPT2 2019 OpenAI Open-source Decoder-only Text generation Fairness and Bias, Robustness
@@
Models Release Year Institution Openness Architecture Primary Task Dimensions
PMC Llama 13B 2023 Allen Institute for AI Open-source Decoder-only Specialized in medical literature understanding and generation. Fairness and Bias, Robustness
@@
GPT-4 2023 OpenAI Closed-source Decoder-only Advanced language generation and understanding across various domains. Fairness and Bias, Safety, Robustness, Explainability, Privacy
@@
BERT 2018 Google AI Language Open-source Encoder-only Pre-trained Transformer model for a wide range of NLP tasks, such as text classification, NER, QA, etc. Fairness and Bias, Safety, Robustness, Truthfulness
@@
LLAMA 2 CHAT 2023 Meta AI Open-source Decoder-only Language modeling Robustness, Explainability
@@
MEDALPACA (7B) 2023 medalpaca Open-source Decoder-only Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Robustness, Privacy
@@
CLINICAL CAMEL (13B) 2023 the AI and healthcare community Open-source Decoder-only Fine-tuned for clinical applications. It is designed to assist with tasks like medical text classification, clinical decision support, information extraction from medical records, and answering clinical questions. Robustness
@@
GPT-2 XL 2019 OpenAI Open-source Decoder-only Large-scale language model for text generation and understanding. Robustness
@@
T5-Large 2020 Google Research Open-source Encoder-decoder It treats all NLP tasks as text-to-text tasks, meaning both the input and output are in the form of text, and it’s used for tasks like translation, summarization, and question answering. Robustness
@@
claude-3.5-sonnet 2024 Anthropic Closed-source Decoder-only It is a variant of Claude, specialized in tasks such as conversational AI, creative writing, poetry generation, and other text-based applications. Robustness
@@
OpenBioLLM-70B 2024 OpenBioAI Open-source Decoder-only It is designed to handle tasks such as biological information extraction, gene sequence analysis, protein folding predictions, and other bioinformatics applications. Robustness
@@
BioMistral-7B 2023 Mistral AI Open-source Decoder-only Focused on biomedical and healthcare-related text. Its tasks include medical question answering, clinical document analysis, and medical text summarization. Robustness
@@
Medllama3-v20 2024 MedAI Labs Open-source Decoder-only Designed to assist in healthcare tasks like clinical reasoning, medical question answering, and patient record analysis. Robustness
@@
Models Release Year Institution Openness Architecture Primary Task Dimensions
ASCLEPIUS (7B) 2023 Asclepius AI Open-source Decoder-only Developed for clinical and medical applications, specializing in tasks like diagnosing medical conditions from symptoms, medical text summarization, and extracting structured information from clinical documents. Robustness, Explainability
ALPACA (7B) 2023 Stanford University Open-source Decoder-only Fine-tuned version of the LLaMA model aimed at providing high-quality responses to questions, with an emphasis on maintaining ethical and accurate conversational capabilities in diverse domains. Robustness
@@
Google’s Bard 2023 Google Closed-source Encoder-decoder Conversational AI tool, focused on providing detailed, accurate, and creative responses to user queries. It can handle a variety of tasks, including web search, content generation, and complex QA. Robustness
@@
Text- Davinci-003 2022 OpenAI Closed-source Decoder-only It is an advanced variant of GPT-3. It is designed for a wide range of natural language understanding and generation tasks, such as answering questions, summarizing text, creative writing, translation, and code generation. Robustness, Truthfulness
@@
LLaMa 2-7B 2023 Meta (formerly Facebook AI Research) Open-source Decoder-only Designed to be a general-purpose AI for a wide range of tasks such as text generation, question answering, and summarization, with specific fine-tuning for medical and technical domains. Robustness, Truthfulness, Privacy
@@
ChatGPT 2022 OpenAI Closed-source Decoder-only Conversational AI Robustness, Truthfulness, Explainability, Privacy
@@
Llama-3.1 2024 Meta AI Open-source Decoder-only Multilingual large language model designed for a variety of natural language processing tasks. Safety, privacy
@@
ClinicalCamel-70b 2023 the AI and healthcare community Open-source Decoder-only Medical language model designed for clinical research applications. Safety, Explainability
@@
Med42-70b 2023 M42 Health Open-source Decoder-only Clinical large language model providing high-quality answers to medical questions. Safety, Explainability
@@
GPT-4o 2024 OpenAI Closed-source Decoder-only Multimodal large language model capable of processing and generating text, audio, and images in real time. Safety, Privacy, Explainability
@@
Mistral 2023 Mistral AI Open-source Decoder-only Language model optimized for code generation and reasoning tasks. Safety, Robustness, Explainability
@@
Meditron (7) (70b) 2023 École Polytechnique Fédérale de Lausanne (EPFL) Open-source Decoder-only Medical language model fine-tuned for clinical decision support and medical reasoning. Safety, Robustness, Explainability
@@
Models Release Year Institution Openness Architecture Primary Task Dimensions
Claude-2.1 2023 Anthropic Closed-source Decoder-only General-purpose language model for a wide range of natural language understanding and generation tasks. Safety, Robustness
GPT-J 2021 EleutherAI Open-source Decoder-only Open-source language model for text generation and understanding. Safety, Robustness
Vicuna 2023 UC Berkeley and Microsoft Research Open-source Decoder-only Conversational AI Safety, Robustness, Truthfulness
@@
Medalpaca-13b 2023 medalpaca Open-source Decoder-only Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Safety, Truthfulness, Privacy
@@
GPT-3 2020 OpenAI Closed-source Decoder-only Natural language understanding and generation Truthfulness, Explainability
@@
ALBERT 2019 Google Research Open-source Encoder-only Lighter version of BERT that reduces parameters for efficiency while maintaining performance. It excels in tasks such as text classification, named entity recognition, and question answering. Truthfulness
@@
RoBERTa 2019 Facebook AI Research Open-source Encoder-only Optimized variant of BERT that removes the Next Sentence Prediction task and trains with more data and for longer periods. It is used for tasks like question answering, sentiment analysis, and text classification. Truthfulness
@@
BlueBERT 2019 NIH and Stanford University Open-source Encoder-only BERT-based model pre-trained on clinical and biomedical text. It is designed for healthcare-related tasks, including clinical text classification, named entity recognition, and medical question answering. Truthfulness
@@
ClinicalBERT 2019 University of Pennsylvania Open-source Encoder-only Variant of BERT fine-tuned on clinical texts, tailored for clinical NLP tasks like named entity recognition, clinical event extraction, and question answering in the medical domain. Truthfulness
@@
TAPAS 2020 Google Research Open-source Encoder-only Designed for answering questions based on tabular data. It is used for tasks like extracting structured information from tables and processing queries in tabular datasets. Truthfulness
@@
LLaMA-2 13B 2023 Meta Open-source Decoder-only Advanced variant of Meta’s LLaMA series, designed for text generation, question answering, summarization, and other NLP tasks. Truthfulness, Explainability, Privacy
@@
MPT 2023 MosaicML Open-source Decoder-only General-purpose LLM for text generation, summarization, language understanding, and reasoning tasks. Fine-tuned for downstream applications such as chatbot development, code generation, and other NLP tasks. Truthfulness
@@
Models Release Year Institution Openness Architecture Primary Task Dimensions
BLIP2 2023 Salesforce Open-source Encoder-decoder Bootstrapping language-image pre-training, designed to bridge vision-language models with large language models for improved visual understanding and generation. Truthfulness
InstructBLIP-7b/13b 2023 Salesforce Open-source Encoder-decoder Visual instruction-tuned versions of BLIP-2, utilizing Vicuna-7B and Vicuna-13B language models, respectively, to enhance vision-language understanding through instruction tuning. Truthfulness
LLaVA1.5-7b/13b 2023 Microsoft Open-source Encoder-decoder Large language and vision assistant models with 7B and 13B parameters, respectively, designed for multimodal tasks by integrating visual information into language models. Truthfulness
mPLUGOwl2 2023 Zhejiang University Open-source Encoder-decoder Multimodal pre-trained language model designed to handle various vision-language tasks, including image captioning and visual question answering. Truthfulness
@@
XrayGPT 2023 University of Toronto Open-source Decoder-only Specialized model for generating radiology reports from chest X-ray images, aiming to assist in medical image interpretation. Truthfulness
@@
MiniGPT4 2023 King Abdullah University of Science and Technology Open-source Decoder-only A lightweight multimodal model designed to align vision and language models efficiently, facilitating tasks like image captioning and visual question answering. Truthfulness
@@
RadFM 2023 Stanford University Open-source Decoder-only Foundation model tailored for radiology, focusing on interpreting medical images and integrating findings with clinical language models. Truthfulness
@@
Alpaca-LoRA 2023 Stanford University Open-source Decoder-only It focuses on achieving good performance in tasks such as question answering and personalized dialogue. Truthfulness
@@
Robin- medical 2023 Robin Health Open-source Decoder-only Fine-tuned for medical applications, including clinical decision support, medical question answering, and health record analysis. Truthfulness
@@
Flan-T5 2021 Google Research Open-source Encoder-decoder Optimized for tasks like question answering, text summarization, and sentence classification, across a variety of domains. Truthfulness, Explainability
@@
BioBERT 2019 Korea University Open-source Encoder-only Biomedical language representation learning, enhancing performance on tasks like named entity recognition, relation extraction, and question answering within the biomedical domain. Truthfulness
@@
Falcon Instruct (7B and 40B) 2023 Technology Innovation Institute (TII), UAE. Open-source Decoder-only Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
@@
Models Release Year Institution Openness Architecture Primary Task Dimensions
Mistral Instruct (7B) 2023 Mistral AI Open-source Decoder-only Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
Falcon 2023 Technology Innovation Institute (TII), UAE. Open-source Decoder-only General-purpose language model optimized for text understanding, generation, question answering, and reasoning tasks. Focused on efficient deployment for industry-scale applications. Truthfulness, Robustness
LLaVA-Med 2024 Microsoft Open-source Encoder-decoder Large language and vision assistant for biomedicine, trained to handle visual instruction tasks in the biomedical field, aiming for capabilities similar to GPT-4. Truthfulness, Explainability
Claude-3 2024 Anthropic Closed-source Decoder-only General-purpose LLM (QA, dialogue, reasoning, summarization) Explainability
GPT-4o-mini 2024 OpenAI Closed-source Decoder-only Natural language processing (NLP), text generation, and understanding. Explainability
ASCLEPIUS (13B) 2023 Asclepius AI Open-source Decoder-only Medical NLP, clinical text analysis, and healthcare-related tasks. Explainability
@@
MedViLaM 2023 Cite Open-source Encoder-decoder Medical vision-language tasks, combining image and text analysis for healthcare. Explainability
@@
Med-MoE 2023 Cite Open-source Decoder-only Medical NLP, leveraging Mixture of Experts (MoE) for specialized healthcare tasks. Explainability
@@
Gemini Pro 2023 Google DeepMind Closed-source Decoder-only Multi-modal NLP, combining text, image, and other data types for advanced AI tasks Explainability
Gemini-1.5 2024 Google DeepMind Closed-source Decoder-only Multimodal reasoning, long-context understanding, QA, generation Explainability
@@
AlpaCare (7B) (13B) 2023 Cite Open-source Decoder-only Healthcare-focused NLP, clinical text analysis, and medical decision support Explainability
@@
Yi (6B) 2023 01.AI (China) Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
Models Release Year Institution Openness Architecture Primary Task Dimensions
@@
Phi-2 (2.7B) 2023 Microsoft Open-source Decoder-only Lightweight NLP, text generation, and fine-tuning for specific tasks. Explainability